Statistics from a​ college's climate center indicate that the city the college is in gets an average of 35.9​" of rain each​ year, with a standard deviation of 4.2​". Assume that a Normal model applies. Complete parts a through d below.

​a) During what percentage of years does this city get more than 42​" of​ rain? The percentage of years with more than 42​" of rain is nothing​%.
Answer #1

Solution :

Given that ,

mean = = 35.9

standard deviation = = 4.2

P(x > 42) = 1 - P(x < 42)

= 1 - P[(x - ) / < (42 - 35.9) / 4.2)

= 1 - P(z < 1.45)

= 1 - 0.9265

= 0.0735

percentage = 7.35%